OpenStudy (anonymous):

The current population of polar bears is estimated at 22,500. What will the population be in three years if it continues declining at 0.3% per year? A. 201 B. 7718 C. 20536 D. 22298

OpenStudy (mertsj):

The polar bear population is not declining.

OpenStudy (ghazi):

if it gets lowered by 0.3% each year then find 0.3 % of 22,500 and multiply it by 3 \[22,500*\frac{ 0.3 }{ 100 }=225*0.3=67.5\] then multiply 67.5 by three =202.5 polar bears wont be there 22,500-203 =22297

OpenStudy (ghazi):

i have taken 202.5 as 203 because you cant have half polar bear, only exception is when half of the polar bear is paralysed or has lost its body :P

OpenStudy (anonymous):

Lol, so its D right? @ghazi

OpenStudy (ghazi):

correct :)

OpenStudy (anonymous):

what about this one For compound interest savings, what are the interest earnings for the lender the first year for $2,300 saved at 13% annual interest compounded annually? A. $199 B. $299 C. $325 D. $330 @ghazi

OpenStudy (ghazi):

sorry its been ages and i have forgot the formula of compound interest lol

OpenStudy (anonymous):

haha its ok

OpenStudy (mertsj):

\[A=P(1+\frac{r}{n})^{nt}\]